# Discovery of Critical Thresholds in Mixed Exposures and Estimation of Policy Intervention Effects using Targeted Learning

### v2

@@ 7.1 NIEHS Synthetic Mixtures
The NIEHS synthetic mixture data (found here in github) are a commonly used data set to evaluate the performance of statistical methods for mixtures. These synthetic data can be considered as the results of a prospective cohort study. The outcome cannot cause exposures (as might occur in a cross-sectional study). Correlations between exposure variables can be thought of as caused by common sources or modes of exposure. The nuisance variable $W$ can be assumed to be a potential confounder and not a collider. There are 7 exposures ($A_{1}-A_{7}$) that have a complicated dependency structure with a biologically based dose response function based on endocrine disruption. For details, the synthetic data key on the github page for data set 1 (used here) gives a description of how the data were generated. Generally, there are two groups of exposures ($A_{1},A_{2},A_{3}$ and $A_{5},A_{6}$). Therefore, the correlations within these clusters are high. $A_{1},A_{2},A_{7}$ contribute positively to the outcome; $A_{4},A_{5}$ contribute negatively; $A_{3}$ and $A_{6}$ do not have an impact on the outcome, making rejecting these variables difficult given their correlations with members of the group. This correlation and effects structure is biologically plausible because different congeners of a group of compounds (e.g., PCBs) may be highly correlated but have different biological effects. There are various agonistic and antagonistic interactions in exposures. A breakdown of the variables sets and their relationships are: $A_{1},A_{2}$ both increase Y by concetration addition, $A_{1}-A_{4},A_{2}-A_{4},A_{1}-A_{4},A_{2}-A_{4}$ competetive antagonism, $A_{1}-A_{7},A_{2}-A_{7}$ supra-additive or synergistic, $A_{4},A_{5}$ both addivively decrease Y, $A_{4}-A_{7},A_{5}-A_{7}$ unusual antagonism. Given these toxicological interactions, we can expect certain statistical interactions determined as cut-points for sets of variables from CVtreeMLE. Since we are interested in one region that minimizes the mean outcome under some restrictions, we expect to find a region that includes lower levels for an exposure with an agonistic relationship or perhaps an interaction between an agonistic and antagonistic exposure. The NIEHS data set has 500 observations and 9 variables. $A$ is a binary confounder. Of course, in this data there is no ground truth, like in the above simulations, but we can gauge CVtreeMLE’s performance by determining if the expected variable sets are included in the rule which defines the minimizing region. We apply CVtreeMLE to this NIEHS synthetic data using a 20-fold CV and the default stacks of estimators used in the Super Learner for all parameters. We parallelize over the cross-validation to test computational run-time on a newer personal machine an analyst might be using. We show results when searching for both the minimizing region and the maximizing region. The minimum number of observations in a region was set to 20.
@@
In 20 of 20 folds, a region in $X_{2}$ was found to have the lowest adjusted outcome. The average rule was found to be $X_{2}<=0.382$. This threshold varied between 0.38 and 0.39 a very low partition variability. The ARE for this region was found to be -4.39 (-6.37 - -2.42), which means that if all individuals were exposed to levels of $X_{2}$ within this region, the result would be 4.47 units less than the observed average. This effect was significant (p < 0.001). The fact that this region was found in 100% of folds indicates a robust result. This result can be interpreted as, in order to have the largest impact to reduce the outcome, this exposure should be adjusted levels below this threshold. For the maximum region with at least 20 observations, we found that X1 has the highest mean outcome at thresholds greater than 3.3 (3.06 - 3.49), restricting observations below this threshold had a borderline significant effect of 8.21 (-3.40 - 19.84). This indicates that if all individuals were forced to this region the outcome would be 8.21 larger than observed average. A threshold for this exposure X1 was found in 100% of folds.
The dose-response curves for these exposures in these synthetic data are provided on the github page. In each case, these results are reasonable given the marginal dose response relationships. Finally, given the antagonistic relationships built in, we might expect that the region with the lowest mean may be an interaction between an agonist and antagonist which we have not detected because of restricting observations in the region, when we reduce the min observations to 10 for the regions, we detect this minimizing region (Table 1). These results give an example of variability, since this minimizing region was found in 90% of the folds that included exposures X2 and X5 and one fold for X2 and X7. The TMLE result of the pooled oracle region was -7.17 (-29.37 - 15.04, p = 0.53) which indicates that the ARE averaged between these two regions is -7.17. Table 1 results provide an example of the results of inverse variance as the detected region changed slightly between folds. From this example, the analyst would report a consistent antagonistic relationship found between exposures X2 and X5 in 90% of the folds with a significant ARE.
ARE
SE
Lower CI
Upper CI
P-value
Average Rule
Prop.
-6.802
1.218
-9.190
-4.413
0.001
X2 <=0.384(0.38,0.39) & X5 >=1.122(0.87,1.23)
0.90
-0.286
6.393
-12.816
12.244
0.96
X2 <= 0.38 & X7 <= 0.44
0.05
Table 1: NIEHS Results: Inverse Variance Estimates for the Minimizing Interaction
@@ 7.2 NHANES Data
Table 3: Inverse Variance Pooled TMLE Results for NHANES POP Exposures
ARE
SD E.
Lower CI
Upper CI
P-value
Average Rule
Prop.
-0.051
0.26
-0.566
0.465
0.85
cadmium $>$ 2.147 (2.04,2.32)
0.30
-0.173
0.965
-2.065
1.719
0.86
cadmium $>$ 1.93 & cesium $>$ 7.4
0.10
-0.154
0.390
-0.919
0.611
0.7
cadmium $\geq$ 1.888 (1.81,2.03) & tungsten $\leq$0.098 (0.09,0.11)
0.60
The ARE for the oracle parameter across the folds was found to be -0.22 (-0.98 - 0.01), p = 0.06. This indicates that for a region in the metal mixture space that minimizes telomere length, if all individuals were exposed to this region the telomere length would reduce by 0.22. Given this omnibus result, we now review Table 3 which shows the inverse pooled TMLE ARE results for the minimal region found in the folds that make up this oracle region. Here, we see that the detected region that minimizes the average telomere length varies slightly across the folds; however, all regions include cadmium. The region that is the most consistent (60 % of folds) is when cadmium is at or above levels of 1.89 (1.81 - 2.03) and tungsten is at or below levels of 0.098 (0.09 - 0.11). If our NHANES population were forced to be exposed to these levels of cadmium and tungsten, we expect 0.154 base pairs to be shorter than the length of the telomere under observed exposures to cadmium and tungsten. This result was found in 6 of 10 of our folds (running a higher CV can make this more stable). The rules presented are for the average partition across the folds; the numbers in brackets show the lower and upper bounds of the partitions for the respective variable in the rule across the folds and thus can be used to assess stability. We simply imputed missing values at the mean; a more thorough review of better preprocessing methods may improve consistency of results, this is simply an applied example.
@@
Our study highlights the need for further research on the health impacts of cadmium and tungsten interactions, particularly on telomere length. Cadmium, a known toxic metal, is linked to health issues like oxidative stress and shorter telomeres. Tungsten, less toxic and less studied, might interact with cadmium in the body, possibly mitigating its harmful effects. The precise nature of this interaction and its effect on cellular aging biomarkers, like telomeres, remains unclear and warrants deeper investigation. This evidence suggests that, when mixed metals are regulated with the aim of reducing the effects on telomere length, the regulation of cadmium and tungsten should be prioritized. Additionally, if populations are exposed to lower levels of tungsten, the regulatory level of cadmium should be lower.
@@ 8 Software
The CVtreeMLE R package, is a pivotal tool for using our proposed method. It offers comprehensive documentation, including a vignette on semiparametric theory, real-life examples, and comparative analyses. Emphasizing reproducibility, the package includes applications with NIEHS synthetic data and NHANES mixed metal exposure data. Notably, it supports efficient computational performance, suitable for standard personal computers and scalable to high-performance computing for simulations. The package, available on GitHub, exemplifies the push towards accessible, well-documented, and robust statistical software in scientific research.
@@ 9 Discussion
In our study, we introduce a novel approach for analyzing mixed exposures, centering on the ’oracle target parameter’ to determine specific exposure regions’ outcomes. This method employs a decision tree and plug-in g-computation for estimations of these regions. Our goal is to efficiently estimate intervention effects on the maximizing or minimizing region in each fold, utilizing targeted learning within a cross-validation framework to estimate the ARE for this region. This research has significant public policy implications, particularly in identifying safe exposure levels and assessing the impacts of various exposure combinations. A limitation is the inverse variance weighting used to pool within exposure set specific regions to aid in interpretability and gain greater power for these estimates rather than relying on k-fold specific results. Future work will focus on more flexible decision tree development and new methods for incorporating null fold findings with testing. We demonstrate the method’s robustness through simulations and real-world data, culminating in the development of the CVtreeMLE R package for wider application.

### v3

@@ 7.1 NIEHS Synthetic Mixtures
The NIEHS synthetic mixture data (found here in github) are a commonly used data set to evaluate the performance of statistical methods for mixtures. These synthetic data can be considered as the results of a prospective cohort study. The outcome cannot cause exposures (as might occur in a cross-sectional study). Correlations between exposure variables can be thought of as caused by common sources or modes of exposure. The nuisance variable $W$ can be assumed to be a potential confounder and not a collider. There are 7 exposures ($A_{1}-A_{7}$) that have a complicated dependency structure with a biologically based dose response function based on endocrine disruption. For details, the synthetic data key on the github page for data set 1 (used here) gives a description of how the data were generated. Generally, there are two groups of exposures ($A_{1},A_{2},A_{3}$ and $A_{5},A_{6}$). Therefore, the correlations within these clusters are high. $A_{1},A_{2},A_{7}$ contribute positively to the outcome; $A_{4},A_{5}$ contribute negatively; $A_{3}$ and $A_{6}$ do not have an impact on the outcome, making rejecting these variables difficult given their correlations with members of the group. This correlation and effects structure is biologically plausible because different congeners of a group of compounds (e.g., PCBs) may be highly correlated but have different biological effects. There are various agonistic and antagonistic interactions in exposures. Table 1 shows the interactions built into this synthetic data.
Variables Interaction Type
A1 and A2 Toxic equivalency factor, a special case of concentration addition (both increase Y)
A1 and A4 Competitive antagonism (similarly for A2 and A4)
A1 and A5 Competitive antagonism (similarly for A2 and A4)
A1 and A7 Supra-additive (“synergy”) (similarly for A2 and A7)
A4 and A5 Toxic equivalency factor, a type of concentration addition (both decrease y)
A4 and A7 Antagonism (unusual kind) (similarly for A5 and A7)
Table 1: NIEHS Synthetic Data Interactions
Given these toxicological interactions, we can expect certain statistical interactions determined as cut-points for sets of variables from CVtreeMLE. Since we are interested in one region that minimizes the mean outcome under some restrictions, we expect to find a region that includes lower levels for an exposure with an agonistic relationship or perhaps an interaction between an agonistic and antagonistic exposure. For example, the synergistic relationship between A1 and A7 may indicate we should decrease these to a certain level simultaneously. The NIEHS data set has 500 observations and 9 variables. $A$ is a binary confounder. We apply CVtreeMLE to this NIEHS synthetic data using a 10-fold CV and the default stacks of estimators used in the Super Learner for all parameters. We parallelize over the cross-validation to test computational run-time on a newer personal machine an analyst might be using. We show results when searching for the minimizing region.
7.2 NIEHS Data Results
In 10 of 10 folds, it was found that a region in $X_{1}\&X7$ had the lowest adjusted outcome. Table 2 shows the k-fold specific results. In this table, the ARE is the change in outcome if all individuals were in the region defined in the column ”region” compared to the observed outcome. Standard errors and confidence intervals are derived from the fold-specific influence function based on the validation data.
Table 2: NIEHS K-fold ARE Results
Fold ARE SE Lower CI Upper CI P-value Adjusted P-value Region
1 -4.33 27.86 -58.94 50.29 0.88 1.00 $X1\leq 0.98\&X7\leq 0.5$
2 -2.88 16.90 -36.00 30.25 0.86 1.00 $X1\leq 1\&X7\leq 0.45$
3 -2.39 20.04 -41.66 36.88 0.91 1.00 $X1\leq 1\&X7\leq 0.46$
4 -3.49 24.92 -52.34 45.35 0.89 1.00 $X1\leq 1\&X7\leq 0.33$
5 -3.88 23.85 -50.63 42.87 0.87 1.00 $X1\leq 1\&X7\leq 0.44$
6 -2.98 21.85 -45.81 39.85 0.89 1.00 $X1\leq 1\&X7\leq 0.52$
7 -4.12 18.73 -40.83 32.58 0.83 1.00 $X1\leq 0.99\&X7\leq 0.35$
8 -4.52 25.85 -55.18 46.14 0.86 1.00 $X1\leq 1\&X7\leq 0.4$
9 -3.98 20.39 -43.95 35.98 0.85 1.00 $X1\leq 1\&X7\leq 0.42$
10 -3.13 27.96 -57.93 51.68 0.91 1.00 $X1\leq 1\&X7\leq 0.4$
Because these K-fold estimates are for very similar regions in $X1$ and $X7$ the pooled TMLE parameter is interpretable. Table 3 shows the pooled results. Here, the “Mixture ARE” column is the average ARE estimates across the folds. The standard error and confidence intervals are now derived from all the validation fold estimates.
Table 3: Summary of mixture average treatment effects (ATE) across different folds.
Fold Mixture ARE Standard Error Lower CI Upper CI P-value Average Rule
Pooled TMLE -3.73 7.31 -18.06 10.59 0.61 $X1\leq 0.997$ $(0.98,1)$ $X7\leq 0.427$ $(0.33,0.52)$
@@ 7.2 NIEHS Data Results
The average rule was found to be $X_{1}<=0.997\&X_{7}<=0.43$. This threshold varied between 0.98 and 1 for $X_{1}$ and 0.33 and 0.52 for $X_{7}$, a very low partition variability. The ARE for this region was found to be -3.73 (–18.06 - 10.59), which means that if all individuals were exposed to levels of $X_{1}$ and $X_{7}$ within this region, the endocrine disrupting outcome would be 3.73 units less than the observed average. This effect was not significant. The fact that this region was found in 100% of folds indicates a robust result. This result can be interpreted as, in order to have the largest impact to reduce the endocrine disrupting outcome, this exposure should be adjusted to levels below this threshold. According to the description of this data found here on github the exposures $X_{1}$ and $X_{7}$ have the strongest synergistic interaction. Therefore, our findings make sense insofar as, to have the greatest reduction in the outcome, both exposures must be reduced simultaneously.
@@ 7.3 NHANES Data
Table 5: Mixtures of Metals and Telomere Length Across Different Folds
Fold ARE SE Lower CI Upper CI p-value Mix Rule
1 -0.54 0.02 -0.58 -0.50 p < 0.001 barium > 0.37 & molybdenum $\leq$ 112.4
2 -0.68 0.02 -0.72 -0.64 p < 0.001 barium > 0.39 & molybdenum $\leq$ 110
3 -0.99 0.02 -1.03 -0.95 p < 0.001 barium > 0.42 & molybdenum $\leq$ 113.9
4 -0.26 0.02 -0.29 -0.22 p < 0.001 barium > 0.41 & barium > 0.63
5 -0.60 0.02 -0.64 -0.56 p < 0.001 barium > 0.41 & molybdenum $\leq$ 109.7
6 -0.69 0.02 -0.74 -0.65 p < 0.001 barium > 0.43 & molybdenum $\leq$ 113.3
7 -2.58 0.05 -2.68 -2.48 p < 0.001 barium > 0.4 & cesium > 2.23
8 -0.96 0.02 -1.00 -0.92 p < 0.001 barium > 0.4 & cesium > 2.1
9 -0.70 0.02 -0.75 -0.66 p < 0.001 barium > 0.42 & molybdenum $\leq$ 119.2
10 -0.71 0.02 -0.75 -0.67 p < 0.001 barium > 0.42 & molybdenum $\leq$ 112
7.4 NHANES Mixed Metal Results
Table 5 shows the k-fold specific results. The results reveal significant associations in each fold with consistent patterns of metal mixtures leading to decreased telomere length. For example, higher levels of barium combined with varying thresholds for lower molybdenum were consistently associated with reduced telomere lengths, as evidenced by consistently negative AREs across several folds. In particular, the combination of barium (>0.42) and molybdenum (<113.9) in Fold 3 showed one of the most substantial decreases in telomere length (ARE = -0.99), with all reported p-values less than 0.001, which confirms the robustness of these findings. This pattern suggests a repeatable dose-response relationship in which specific levels of barium and molybdenum contribute to the shortening of the telomere. In one fold, fold 7, the minimizing region was in barium and cesium and in one fold, fold 4, the minimizing region was for barium alone without interaction. The fact that the minimizing region was found for high levels of barium (>0.4) and low levels of molybdenum < 113 in 7 of 10 folds indicates that these two chemicals have an antagonistic relationship which in combination can lead to the most severe reduction in telomere length. Running in more folds (such as 20) could make this result more consistent.
Table 6: Pooled Estimation of Regional Attributable Effect
Fold Region ARE Standard Error Lower CI Upper CI P-value
Pooled TMLE -2.52 0.01 -2.54 -2.50 $<$0.001
The ARE for the oracle parameter across the folds was found to be -2.52 (-2.54 - -2.50), p < 0.001. This indicates that for a region in the metal mixture space that minimizes telomere length, if all individuals were exposed to this region the telomere length would reduce by 2.52. Of course, as seen above in the k-fold specific results, this region has some variability and therefore is more difficult to interpret. It is interpreted as, telomere length reduces by 2.52 if all individuals were exposed to the estimated most severe combination of metals.
@@ 7.4 NHANES Mixed Metal Results
Individual studies have demonstrated that prenatal exposure to specific metals such as barium and molybdenum may impact telomere length in newborns. Cowell et al. [2020] found a significant association where barium contributed to the strongest inverse relationship with newborn telomere length, suggesting that exposure to this metal may accelerate cellular aging processes as indicated by shortened telomeres. Complementarily, Xia et al. [2022] reported a positive association between urinary molybdenum and telomere length, particularly in the elderly, which might suggest a protective or mitigating effect of this metal on telomere attrition. These findings are consistent with our results using the CVtreeMLE approach, which not only reaffirms the roles of these metals when analyzed individually but also enhances our understanding of their combined effects on telomere biology. This multi-metal interaction analysis provides a more nuanced view of how exposure to a mixture of metals like barium and molybdenum may influence telomere length and, by extension, long-term health outcomes.
@@ 8 Software
The CVtreeMLE R package, implements our proposed methodology. It offers comprehensive documentation, including a vignette on semiparametric theory, real-life examples, and comparative analyses. Emphasizing reproducibility, the package includes applications with NIEHS synthetic data and NHANES mixed metal exposure data. Notably, it supports efficient computational performance, suitable for standard personal computers and scalable to high-performance computing for simulations. The package, available on GitHub, exemplifies the push towards accessible, well-documented, and robust statistical software in scientific research.
@@ 9 Discussion
In this study, we present a novel approach for analyzing mixed exposures, focusing on the data-adaptive identification of critical thresholds in multivariate exposure spaces. Our method introduces a decision tree algorithm integrated with targeted learning to estimate intervention effects within identified exposure regions. This approach leverages cross-validated targeted maximum likelihood estimation to ensure robust and efficient estimation of the attributable regional effect for these regions, making it particularly suitable for complex environmental health data.
[…]
Through extensive simulations and application to real-world data, including NIEHS synthetic mixtures and NHANES metal exposure data, we demonstrate the robustness and practical applicability of our method. The simulation studies confirm the asymptotic convergence of our estimator to the true oracle region and its effectiveness in identifying the correct thresholds. The real-world applications further illustrate how our method can identify harmful exposure levels and provide actionable insights for public health interventions.
However, our approach has limitations. When the identified regions differ significantly across folds, it indicates weak signal strength, making the interpretation of the pooled parameter challenging. This variability suggests that the data do not strongly support a single optimal exposure region, complicating regulatory decisions. In such cases, environmental regulators must carefully assess the fold-specific results to understand the potential exposure thresholds better. This limitation underscores the importance of considering the consistency and stability of the identified regions across different data subsets.
One key aspect of our methodology is the preservation of relative self-selection preferences within the identified regions. This choice aligns with existing literature that often binarizes continuous exposures to facilitate causal inference. By maintaining the natural selection preferences within exposure regions, our approach aims to provide familiar and interpretable estimates of the impact of policy interventions. However, the self-selection strategy may not always represent the most plausible policy intervention. For instance, a policy that moves individuals just below a threshold could result in higher densities near the boundary and lower densities farther from the threshold, reflecting more realistic enforcement scenarios. Expanding our approach to estimate different policies based on domain knowledge and specific regulatory contexts will enhance its applicability and relevance. Future research should explore methods to incorporate such domain-specific policies, providing a more comprehensive toolkit for environmental regulators and policymakers.
Despite these limitations, our method represents a significant advancement in the analysis of mixed exposures. It offers a flexible and powerful tool for uncovering critical exposure thresholds and assessing the combined effects of multiple environmental exposures. The development of the CVtreeMLE R package facilitates the broader application of our methodology, providing researchers and policymakers with a robust, user-friendly tool for evaluating complex exposure data.
In conclusion, our study provides a comprehensive framework for the data-adaptive identification and estimation of safe exposure thresholds in multivariate settings. This approach has substantial implications for public policy, particularly in setting regulatory limits and developing interventions to mitigate the adverse health effects of environmental exposures. Future research should focus on further refining this method, exploring its application to other types of exposure data, and addressing the challenges related to signal variability across folds.